OpenStudy (anonymous):

3*e^(4-3x)+1.6=45.6 need help solving

OpenStudy (anonymous):

I managed to get e^(4-3x)=44/3 by subtracting 1.6, then dividing 3 from both sides.

OpenStudy (perl):

looks good so far

OpenStudy (anonymous):

Sorry I'm blanking out on what to do next,

OpenStudy (anonymous):

Could you please help me?

OpenStudy (perl):

we can take log of both sides

OpenStudy (perl):

'ln' it

Directrix (directrix):

ln( e ^(4-3x) ) = ln (44/3) where ln is the natural (base e) log.

OpenStudy (anonymous):

well I understand your mail, I need to simplify, and not get any exact answers with decimals, but only in the simplest form. Sorry

OpenStudy (anonymous):

(4-3x)ln(e)=ln(44/33) is this right?

OpenStudy (anonymous):

wht can I do to fix the ln of both side for x?

OpenStudy (anonymous):

sorry I'm real lost

Directrix (directrix):

Yes. (4-3x)ln(e)=ln(44/33) ln(e)= 1 because e^1 =1 So, (4-3x)ln(e)=ln(44/33) becomes 4 - 3x = ln(44/33) So what is the natural log of 44/33? Use your calculator or one online. Then, we will be ready for the next step.

Directrix (directrix):

The Google calculator is helpful: Take a look. https://www.google.com/?gws_rd=ssl#q=natural+log+of+44%2F33

OpenStudy (anonymous):

Sorry I don't think I'm supposed to use a calculator. I think I was supposed to use e and log() and ln(). Is there a way to simplify ln(44/3) into a non-ln() way? If there isn't, i think thats the closest I can get to keep my answers exact.

OpenStudy (perl):

you can keep it as ln(44/3), since there is no way to simplify that in a non-ln( ) way Wait till the end to approximate

Directrix (directrix):

44/3 = 11 so I guess we are looking for the exact natural log of 11. It would be helpful to know the exact instructions for this problem.

Directrix (directrix):

@perl The OP wants exact answers. I don't know if that is altogether possible in this context.

OpenStudy (anonymous):

4-3x=ln(44/3) -3x=ln(44/3)-4 x=-ln(33/4)-4/-3 is this aout right?

OpenStudy (perl):

x = [ ln(44/3) - 4 } / -3 , that is correct

OpenStudy (anonymous):

Would I get full credit if I showed the work? I mean, there is no way to simplify it more, right?

OpenStudy (perl):

there is no way to simplify it more , unless you want to distribute the negative sign

OpenStudy (perl):

x = ( 4 - ln(44/3) ) / 3

OpenStudy (anonymous):

I see, So then it's simplified. Thanks!

OpenStudy (perl):

:)